Circular No. 5606

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1992aw IN ANONYMOUS GALAXY J. Mueller reports her discovery of a supernova 13′′ east and 7′′ north of

the nucleus of an anonymous galaxy at α = 19^{h}05^{m}39^{s}.39, δ = +50^{o}58′05′′.9 (equinox 1950.0). It was of red magnitude ∼ 18 on a plate taken on Aug. 27 with the 1.2-m Oschin Telescope in the course of the second Palomar Sky Survey. A spectrum obtained by S. M. Hughes, J. R. Mould and M. Rich with the double spectrograph on the 5-m Hale telescope on Sept. 5 shows the object to be of type II. Preliminary measurement of the redshift of the galaxy and the FWHM of the Hα emission line in the supernova gives z = 0.036 and 8000 km/s.”

GRO J0422+32 C.-Y. Shao, Harvard-Smithsonian Center for Astrophysics, reports:

“Over 300 patrol plates from the Harvard plate collection were examined to detect any previous outburst of the optical counterpart of this source. The plates give fairly complete monthly coverage of the region during 1928-1953 and during 1967-1989, but no image down to mag 14-15 was found at the position. Sporadic, larger-scale plates were taken between 1910 and 1940, and these show nothing down to mag 16-17. An exposure with the Oak Ridge Observatory’s 0.40-m astrograph on 1992 Sept. 2.3 UT shows the star at m_{pg} = 13.8.”

NOTICE TO CONTRIBUTORS Contributors are advised of the care needed in giving the correct equi-

nox for the positions of the objects they observe. Following a resolution by IAU Commission 20, positions of comets and minor planets are now to be given in the FK_{5}/J2000.0 system; if the position reported is still in terms of 1950.0, this should be specifically indicated. There have not been similar policy changes by the IAU commissions concerned with novae and supernovae, and observers of such objects should therefore generally continue to use the 1950.0 equinox. Other discoveries, particularly those made at other than optical wavelengths, frequently receive designations that depend on position. In such cases the default position is assumed to be 1950.0. If the 2000.0 position is used instead, that part of the designation indicating the position should be immediately preceded by the letter “J”. The designation RE 0457–28 given on IAUC 5518 is thus incorrect; it should have been given as RE J0457–28.

1992 September 5 Brian G. Marsden